one mole of ideal gas (avogadro number: 6.02*10^23 molecules) occupies the 22.4 litres under normal condition. What fraction of this volume is actually occupied by the ideal gas molecules? Treat the molecules of gas as sphere of radius of 1 angstroms.

state your answer to the nearest 0.00001


1
Expert's answer
2021-09-30T14:24:22-0400

The volume of one molecule


"V_0=\frac{4}{3}\pi R^3\\=\frac{4}{3}*3.14*(1*10^{-10}\rm \:m)^3=4.19*10^{-30}\:m^3"

So


"\frac{V_N}{V}=\frac{NV_0}{V}\\=\frac{6.02*10^{23}*4.19*10^{-30}\:\rm m^3}{22.4*10^{-3}\:\rm m^3}=0.00011"